Starting with the data values 70 and 100, add three data values to the sample so that the mean is 92, the median is 91, and the mode is 91.

Answer by vksarvepalli(154)

let the data values to be added be a, b & c.
for mean to be 92
(70+100+a+b+c)/5=92
for mode to be 91 among the five values the maximum occurring value should be 91
so there should be either three or atleast two 91's among the five
if all a,b, and c are = 91 then their mean will not be = 92
so two are = 91
let a= 91 and b=91
now (70+91+91+100+c)/5=92
=> 352+c=460
so c=108
now we will get that the values will be [ 70 91 91 100 108 ] Ans
also the median the center one when arranged in asc or dec order is 91 as required.